OpenStudy (anonymous):

d/dx * [(integral of (1-t^2) dt from 3 to sin x)]....ill write it nicer with the equation box

OpenStudy (anonymous):

\[d/dx ( \int\limits_{3}^{sinx} 1-t ^{2}) dt ) \]

OpenStudy (zarkon):

\[\frac{d}{dx}\int\limits_{a}^{g(x)}f(t)dt=f(g(x))g'(x)\]

OpenStudy (anonymous):

im not sure how to substitute it into taht equation

OpenStudy (zarkon):

\[f(t)=1-t^2\] \[g(x)=\sin(x)\] \[a=3\]

OpenStudy (zarkon):

does that help?

OpenStudy (anonymous):

yes so now it becomes f of sinx*cosx?

OpenStudy (zarkon):

\[g'(x)=\cos(x)\] \[f(g(x))=1-\sin^2(x)=\cos^2(x)\] \[f(g(x))g'(x)=\cos^3(x)\]

OpenStudy (anonymous):

so that is the final answer?

OpenStudy (zarkon):

what do you think?

OpenStudy (anonymous):

yup

OpenStudy (zarkon):

it is

OpenStudy (anonymous):

oh cool..thnks for your help! u explain very well